log4(x-2)+log4(x+10)=3
Answer by stanbon(75887)

log4(x-2)+log4(x+10)=3
------
log4[(x-2)(x+10)] = 3
------
log4(x^2+8x-20) = 3
------
x^2+8x-20 = 4^3 = 64
-----
x^2 + 8x - 84 = 0
(x-6)(x+14) = 0
----
Positive solution:
x = 6
=================
Cheers,
Stan H.
=================